# thanks: Corresponding author

### v2

@@ Abstract
Real-time and accurate prediction of aerodynamic flow fields around airfoils is crucial for flow control and aerodynamic optimization. However, achieving this remains challenging due to the high computational costs and the non-linear nature of flow physics. Traditional Computational Fluid Dynamics (CFD) methods face limitations in balancing computational efficiency and accuracy, hindering their application in real-time scenarios. To address these challenges, this study presents AeroDiT, a novel surrogate model that integrates scalable diffusion models with transformer architectures to address these challenges. Trained on Reynolds-Averaged Navier-Stokes (RANS) simulation data for high Reynolds-number airfoil flows, AeroDiT accurately captures complex flow patterns while enabling real-time predictions. The model demonstrates impressive performance, with average relative $L_{2}$ errors of 0.1, 0.025, and 0.050 for pressure $p$ and velocity components $u_{x},u_{y}$, confirming its reliability. The transformer-based structure allows for real-time predictions within seconds, outperforming traditional U-net diffusion models. This work underscores the potential of generative machine learning techniques to advance computational fluid dynamics, offering potential solutions to persistent challenges in simulating high-fidelity aerodynamic flows.
@@ I Introduction
In recent years, the emergence of deep learning techniques has opened new avenues for efficient flow field predictions. A growing body of research has integrated deep learning methods with fluid flow simulation, yielding promising results. Notably, neural network architectures such as U-net Guo, Li, and Iorio (2016); Sekar et al. (2019); Ribeiro et al. (2020); Chen et al. (2021); Cai et al. (2022), generative adversarial networks (GANs) Goodfellow et al. (2020); Chen, Chiu, and Fuge (2020); Haizhou et al. (2022), and physics-informed neural networks (PINNs) Raissi, Perdikaris, and Karniadakis (2019); Cai et al. (2021); Chen, Liu, and Sun (2021); Zhang et al. (2023) have shown considerable potential in tasks such as flow field reconstruction. These approaches demonstrate the power of deep learning to capture the complex dynamics of fluid flow, signaling a shift in the future of CFD research. Guo et al.Guo, Li, and Iorio (2016) proposed a convolutional neural network (CNN)-based surrogate model for real-time prediction of steady laminar flow fields, achieving significantly faster velocity field estimation compared to traditional CFD solvers while maintaining low error rates. Raissi et al.Raissi, Perdikaris, and Karniadakis (2019) introduced physics-informed neural networks , a deep learning framework that integrates physical laws governed by nonlinear partial differential equations to solve forward and inverse problems, enabling data-efficient spatio-temporal function approximation and scientific discovery even in small data regimes.
The deep learning landscape is still rapidly evolving, with newer architectures—such as Transformers Vaswani (2017); Khan et al. (2022); Lin et al. (2022) and denoising diffusion probabilistic models (DDPMs) Ho, Jain, and Abbeel (2020); Yang and Sommer (2023, 2023)—along with scalable diffusion models integrated with Transformers (DiT) Peebles and Xie (2023), showing superior performance in tasks like image generation. This evolution raises an important question: Can the latest advancements in machine learning enable real-time, accurate predictions of flow fields, effectively addressing the challenges of computational efficiency and accuracy?
@@
While these works mentioned above collectively aim to replace high-cost simulation iterations via a paradigm integrating physical constraints with data-driven intelligence, they exhibit limitations: conventional generative models (e.g., GANs, DDPMs) inadequately capture complex separated flows in high-Reynolds-number turbulence; U-Net-based diffusion models suffer from inefficient iterative sampling that precludes real-time application; and purely data-driven approaches show weak extrapolation capability under extreme Reynolds numbers while lacking embedded physical constraints, leading to localized errors near critical airfoil surfaces. However, due to the complexity of the flow—characterized by turbulence and flow separation—achieving both efficiency and accuracy in these predictions remains a significant challenge.
This study introduces a novel methodology for aerodynamic flow-field simulation—Adaptive Diffusion Transformers for Airfoil Flow Simulation (AeroDiT)—which employs diffusion transformers to learn from high-fidelity Reynolds-Averaged Navier–Stokes (RANS) simulation data. By treating the complex flow-field distributions as high-dimensional probability spaces and leveraging a generative approach, AeroDiT aims to dramatically reduce computational costs while preserving predictive accuracy at high Reynolds numbers. Compared to conventional computational fluid dynamics (CFD) techniques, AeroDiT’s innovations extend both the theoretical and practical frontiers of aerodynamic simulation, yielding the following key contributions:
@@
2. Experimentally Validated, High-Fidelity, Real-Time Prediction Capability. During validation, AeroDiT exhibits exceptional accuracy and real-time performance. Across a variety of angles of attack, freestream velocities, and representative airfoil geometries, the model consistently achieves relative L2 prediction errors on velocity fields, pressure distributions, and turbulence metrics (such as vorticity) on the order of only a few percent. Simultaneously, the generative inference process can reconstruct a complete flow field in a matter of seconds—many orders of magnitude faster than RANS-based CFD. Consequently, AeroDiT holds immense promise for applications requiring massive parametric sweeps, real-time control, or rapid design iteration. It can accelerate the design optimization cycle in aerospace engineering and furnish reliable surrogate models for data-driven, multidisciplinary design optimization.
@@
The remainder of the paper is organized as follows: Section II reviews related work our work has been built on, and provide an overview of the background information. Section III outlines the fundamental methodology used in this work. Section IV presents the experimental evaluation, detailing datasets, configurations, and results obtained through extensive testing. Section V discusses the implications of the results, highlighting the strengths and limitations of the proposed approach while suggesting future research directions.
@@ II.1 Diffusion Models in Fluid Dynamics
Recent advances in deep learning-based fluid dynamics prediction have utilized generative models to address the challenges of simulating complex flow fields. Yang et al. Yang and Sommer (2023) introduce FluidDiff, a novel generative model based on Denoising Diffusion Probabilistic Models (DDPMs). FluidDiff employs a diffusion process to learn a high-dimensional representation of dynamic systems and uses Langevin sampling to predict flow states under specific initial conditions. Trained on finite, discrete fluid simulation data, FluidDiff achieves competitive performance, accurately predicting test data without the need for explicit encoding of the underlying physical system’s prior knowledge. Similarly, Shu et al. Shu, Li, and Farimani (2023) propose a DDPM-based model for flow field super-resolution, trained on high-fidelity data to reconstruct detailed flow fields from low-fidelity or sparsely sampled data. This model also incorporates physics-informed conditioning through partial differential equations (PDEs), improving its predictive accuracy. Hu et al. Hu, Lu, and Yang (2024) introduce the Geometry-to-Flow (G2F) diffusion model, which generates flow fields around obstacles by conditioning on obstacle geometry. Zhou et al. Zhou et al. (2024) present Text2PDE, a latent diffusion model framework for PDE simulation that uses autoencoders and text-based conditioning to generate full spatio-temporal solutions. Liu et al.Liu and Thuerey (2024) proposed an uncertainty-aware surrogate model for turbulence simulations using DDPMs, demonstrating superior accuracy and uncertainty quantification compared to Bayesian neural networks and heteroscedastic models. While these studies show promising results, many of them primarily focus on simpler flow fields, leaving room for further exploration in more complex and high-dimensional flow prediction scenarios. To address these limitations, there is increasing interest in developing more advanced architectures capable of modeling such complexities, such as the Diffusion Transformer, which is introduced in the next section.
@@ III Methodology
Figure 1: The framework and learning process of the proposed AeroDiT (Normalized flow field of EPPLER 59 airfoil).
This study employs Diffusion Transformers to predict the airfoil flow field at high Reynolds numbers. The framework of Diffusion Transformers (DiTs) is illustrated in Figure 1. By leveraging the generative capabilities of diffusion models, DiTs function as a surrogate model that effectively learns the underlying statistical distribution of flow field data. The model is designed to predict the flow characteristics, such as velocity components and pressure fields, based on input features including the geometry of the airfoil and various flow conditions. The overall architecture consists of three primary components: the Condition Encoder, the Diffusion Transformer (DiT) block, and the Function Space Decoder. The Condition Encoder processes input data such as the airfoil geometry and flow velocities to generate an initial latent representation. This encoded information is then fed into the DiT block, where noise is added, and a series of linear transformations, reshaping, layer normalization, and attention mechanisms take place to effectively model the underlying data distribution. The DiT block utilizes multi-head self-attention, pointwise feed-forward networks, and residual connections to capture the complex relationships between the variables in the latent space. Finally, the Function Space Decoder decodes the learned representation into meaningful output data, such as the predicted pressure field and velocity components (both horizontal and parallel), which are visualized in the final output.
Once trained, DiTs can be used as a predictive tool, providing accurate approximations of the flow field for different aerodynamic configurations, without the need for traditional computational fluid dynamics (CFD) simulations. This offers a computationally efficient and scalable alternative for real-time flow field predictions based on airfoil geometry and specified flow conditions. The geometric shape of the wing is shown in Figure 2, with blue for the flow field region and red for the wing shape.
The training process consists of two key steps: forward diffusion and reverse denoising, as shown in Figure 3. In the forward diffusion step, Gaussian noise is progressively added to the input flow field, gradually transforming it into a noise distribution. The reverse denoising process then reconstructs the original flow field iteratively, utilizing Bayesian inference to model the transition probabilities. Together, these processes enable the model to effectively approximate the underlying distribution and generate accurate predictions. The following sections introduce the relevant notations and learning objectives, followed by a detailed description of the forward and reverse diffusion processes.
@@ III.2 Learning Target
The primary objective is to develop a surrogate model, $\mathbf{S}$, that approximates the real physical system, $\mathbf{R}$, governed by the Navier-Stokes equations. As discussed before, solving these equations numerically for real flow fields $\mathbf{R}(\bm{\Psi})$ is computationally expensive. To address this, the learning target is defined as identifying the optimal parameters $\theta$ of the surrogate model $\mathbf{S}$ such that:
@@
III.3 Forward Diffusion Process
@@
III.4 Reverse Denoising Process and the Loss
@@
III.5 Overall Training
As shown in Figure 1, the model is trained in a sequential manner, starting with the encoder, followed by the decoder, and finally the DiT network. Specifically, an asymmetric training approach is employed for the conditional encoder and the functional space decoder.
The conditional encoder aims to extract task-relevant features from the input conditional data. By incorporating specific conditional information, it generates conditionally enriched feature representations, which are then fed into the DiT network to guide the generation of conditional output features in the diffusion process.
The functional space decoder plays a pivotal role in mapping the features produced by the DiT network to the target task space, effectively achieving the specific task objectives for data generation. During feature decoding, it leverages task-specific properties to enhance the model’s generalization performance.
To train DiT module, we employ the MSE loss $\mathcal{L}_{MSE}(\theta)$ to quantify the discrepancy between the noise predicted by DiT module and the true noise. The DiT training procedures are summarized in Algorithm 1.

### v3

@@ Abstract
Real-time and accurate prediction of aerodynamic flow fields around airfoils is crucial for flow control and aerodynamic optimization. However, achieving this remains challenging due to the high computational costs and the non-linear nature of flow physics. Traditional Computational Fluid Dynamics (CFD) methods face limitations in balancing computational efficiency and accuracy, hindering their application in real-time scenarios. To address these challenges, this study presents AeroDiT, a novel surrogate model that integrates scalable diffusion models with transformer architectures to address these challenges. Trained on Reynolds-Averaged Navier-Stokes (RANS) simulation data for high Reynolds-number airfoil flows, AeroDiT accurately captures complex flow patterns while enabling real-time predictions. The model demonstrates impressive performance, with mean relative $L_{2}$ errors of 0.1, 0.025, and 0.050 for pressure $p$ and velocity components $u_{x},u_{y}$, confirming its reliability. To further enhance physical consistency, we incorporate explicit physics-informed losses based on RANS residuals, including mass and momentum conservation constraints. The transformer-based structure allows for real-time predictions within seconds, enabling efficient aerodynamic simulations. This work underscores the potential of generative machine learning techniques to advance computational fluid dynamics, offering potential solutions to challenges in simulating high-fidelity aerodynamic flows.
@@ I Introduction
In recent years, the emergence of deep learning techniques has opened new avenues for efficient flow field predictions. A growing body of research has integrated deep learning methods with fluid flow simulation, yielding promising results. Notably, neural network architectures such as U-net Guo, Li, and Iorio (2016); Sekar et al. (2019); Ribeiro et al. (2020); Chen et al. (2021); Cai et al. (2022), generative adversarial networks (GANs) Goodfellow et al. (2020); Chen, Chiu, and Fuge (2020); Haizhou et al. (2022), and physics-informed neural networks (PINNs) Raissi, Perdikaris, and Karniadakis (2019); Cai et al. (2021); Chen, Liu, and Sun (2021); Zhang et al. (2023) have shown considerable potential in tasks such as flow field reconstruction. These approaches demonstrate the power of deep learning to capture the complex dynamics of fluid flow, signaling a shift in the future of CFD research. Guo et al.Guo, Li, and Iorio (2016) proposed a convolutional neural network (CNN)-based surrogate model for real-time prediction of steady laminar flow fields, achieving significantly faster velocity field estimation compared to traditional CFD solvers while maintaining low error rates. Wu et al.Haizhou et al. (2022) proposed a data-augmented Generative Adversarial Network (daGAN) for rapid and accurate airfoil flow field prediction, demonstrating strong generalization capabilities with sparse training data through pre-training and fine-tuning modules. Raissi et al.Raissi, Perdikaris, and Karniadakis (2019) introduced physics-informed neural networks , a deep learning framework that integrates physical laws governed by nonlinear partial differential equations to solve forward and inverse problems, enabling data-efficient spatio-temporal function approximation and scientific discovery even in small data regimes.
The deep learning landscape is still rapidly evolving, with newer architectures such as Transformers Vaswani (2017); Khan et al. (2022); Lin et al. (2022) and denoising diffusion probabilistic models (DDPMs) Ho, Jain, and Abbeel (2020); Yang and Sommer (2023, 2023)-along with scalable diffusion models integrated with Transformers (DiT) Peebles and Xie (2023), showing superior performance in tasks like image generation. This evolution raises an important question: Can the latest advancements in machine learning enable real-time, accurate predictions of flow fields, effectively addressing the challenges of computational efficiency and accuracy?
@@
While these works mentioned above collectively aim to replace high-cost simulation iterations via a paradigm integrating physical constraints with data-driven intelligence, they exhibit limitations: conventional generative models (e.g., GANs, DDPMs) inadequately capture complex separated flows in high-Reynolds-number turbulence; U-Net based diffusion models suffer from inefficient iterative sampling that precludes real-time application; and purely data-driven approaches show weak extrapolation capability under extreme Reynolds numbers while lacking embedded physical constraints, leading to localized errors near critical airfoil surfaces. However, due to the complexity of the flow-characterized by turbulence and flow separation-achieving both efficiency and accuracy in these predictions remains a significant challenge.
This study introduces a novel methodology for aerodynamic flow-field simulation Adaptive Diffusion Transformers for Airfoil Flow Simulation (AeroDiT) which employs diffusion transformers to learn from high-fidelity Reynolds-Averaged Navier-Stokes (RANS) simulation data. By treating the complex flow-field distributions as high-dimensional probability spaces and leveraging a generative approach, AeroDiT aims to dramatically reduce computational costs while preserving predictive accuracy at high Reynolds numbers. Compared to conventional computational fluid dynamics (CFD) techniques, AeroDiT’s innovations extend both the theoretical and practical frontiers of aerodynamic simulation, yielding the following key contributions:
@@
2. Experimentally Validated, High-Fidelity, Real-Time Prediction Capability. During validation, AeroDiT exhibits exceptional accuracy and real-time performance. Across a variety of angles of attack, freestream velocities, and representative airfoil geometries, the model consistently achieves relative L2 prediction errors on velocity fields, pressure distributions, and turbulence metrics (such as vorticity) on the order of only a few percent. Simultaneously, the generative inference process can reconstruct a complete flow field in a matter of seconds-many orders of magnitude faster than RANS-based CFD. Consequently, AeroDiT holds immense promise for applications requiring massive parametric sweeps, real-time control, or rapid design iteration. It can accelerate the design optimization cycle in aerospace engineering and furnish reliable surrogate models for data-driven, multidisciplinary design optimization.
@@
The remainder of the paper is organized as follows: Section II reviews related work our work has been built on, and provides an overview of the background information. Section III outlines the fundamental methodology used in this work. Section IV presents the experimental evaluation, detailing datasets, configurations, and results obtained through extensive testing. Section V discusses the implications of the results, highlighting the strengths and limitations of the proposed approach while suggesting future research directions.
@@ II.1 Diffusion Models in Fluid Dynamics
Recent advances in deep learning-based fluid dynamics prediction have utilized generative models to address the challenges of simulating complex flow fields. Yang et al. Yang and Sommer (2023) introduce FluidDiff, a novel generative model based on Denoising Diffusion Probabilistic Models (DDPMs). FluidDiff employs a diffusion process to learn a high-dimensional representation of dynamic systems and uses Langevin sampling to predict flow states under specific initial conditions. Trained on finite, discrete fluid simulation data, FluidDiff achieves competitive performance, accurately predicting test data without the need for explicit encoding of the underlying physical system’ prior knowledge. Similarly, Shu et al. Shu, Li, and Farimani (2023) propose a DDPM-based model for flow field super-resolution, trained on high-fidelity data to reconstruct detailed flow fields from low-fidelity or sparsely sampled data. This model also incorporates physics-informed conditioning through partial differential equations (PDEs), improving its predictive accuracy. Hu et al. Hu, Lu, and Yang (2024) introduce the Geometry-to-Flow (G2F) diffusion model, which generates flow fields around obstacles by conditioning on obstacle geometry. Zhou et al. Zhou et al. (2024) present Text2PDE, a latent diffusion model framework for PDE simulation that uses autoencoders and text-based conditioning to generate full spatio-temporal solutions. Liu et al.Liu and Thuerey (2024) proposed an uncertainty-aware surrogate model for turbulence simulations using DDPMs, demonstrating superior accuracy and uncertainty quantification compared to Bayesian neural networks and heteroscedastic models. While these studies show promising results, many of them primarily focus on simpler flow fields, leaving room for further exploration in more complex and high-dimensional flow prediction scenarios. To address these limitations, there is increasing interest in developing more advanced architectures capable of modeling such complexities, such as the Diffusion Transformer, which is introduced in the next section.
@@ III Methodology
Figure 1: Overview of the AeroDiT framework. The input consists of three channels: airfoil geometry mask, horizontal velocity ($u_{x}$), and vertical velocity ($u_{y}$), concatenated as a tensor of shape $[B,3,H,W]$. These are processed by the Condition Encoder, which downsamples the input into a latent representation. Simultaneously, the ground truth pressure and velocity fields are encoded into a target latent space by a VAE. The Diffusion Transformer (DiT) operates on this latent space by iteratively denoising the noisy latent using transformer blocks conditioned on the encoded input and timestep embeddings. The output latent is decoded by the Function Space Decoder to reconstruct the pressure and velocity fields. The model is optimized using a combination of data loss, mass conservation loss, and momentum conservation loss to ensure physical consistency.
This study employs Diffusion Transformers to predict the airfoil flow field at high Reynolds numbers. The framework of DiTs is shown in Figure 1. By leveraging the generative nature of diffusion models, DiTs serve as a surrogate model that learns the statistical distribution of flow field data. Once trained, the model can predict velocity and pressure fields based on input airfoil geometry and flow conditions.
The training process consists of two key steps: forward diffusion and reverse denoising, as shown in Figure 1. In the forward diffusion step, Gaussian noise is progressively added to the input flow field, gradually transforming it into a noise distribution. The reverse denoising process then reconstructs the original flow field iteratively, utilizing Bayesian inference to model the transition probabilities. Together, these processes enable the model to effectively approximate the underlying distribution and generate accurate predictions.
We then introduce the notations, model architecture, diffusion processes, loss functions, and training procedure employed in AeroDiT.
@@ III.2 Learning Target
The primary objective is to develop a surrogate model, $\mathbf{S}$, that approximates the real physical system, $\mathbf{R}$, governed by the Navier-Stokes equations.
As discussed before, solving these equations numerically for real flow fields $\mathbf{R}(\bm{\Psi})$ is computationally expensive. To address this, the learning target is defined as identifying the optimal parameters $\theta$ of the surrogate model $\mathbf{S}$ such that:
@@
III.3 Model Architecture
Figure 1 illustrates the overall architecture of AeroDiT, which consists of three main modules: a condition encoder, a transformer-based diffusion module (DiT), and a function space decoder. The model operates entirely in the latent space, enabling compact yet expressive representations of high-resolution flow fields.
Input tensor structure. The input to the model is a tensor of shape $[B,3,H,W]$, where $B$ is the batch size, and $H\times W$ denotes the spatial resolution of the flow field (typically $128\times 128$). The three input channels correspond to: (1) the airfoil geometry mask (binary-valued), as illustrated in Figure 3, (2) inflow velocity in the $x$-direction ($u_{x}$), and (3) inflow velocity in the $y$-direction ($u_{y}$).
@@ III.3 Model Architecture
Instead of directly feeding raw spatial inputs into the diffusion model, we first encode them using a dedicated condition encoder, composed of four convolutional layers with stride 2 and ReLU activations. This module downscales the input to a compact latent condition embedding of shape $B,C,H^{\prime},W^{\prime}]$, where $C$ is the latent channel dimension and $H^{\prime},W^{\prime}=H/8,W/8$. This design reduces memory overhead and enhances generalization by extracting task-relevant features while suppressing irrelevant artifacts.
Target VAE encoder. In parallel, the ground truth outputs-pressure $p$, and velocity components $u_{x}$, $u_{y}$-are passed through a variational autoencoder (VAE) encoder of identical structure to obtain a latent representation of the same shape $B,C,H^{\prime},W^{\prime}]$. The total number of parameters in the encoder-decoder pair is approximately 1 million, offering a compact and efficient representation of high-fidelity simulation data. While the encoder is trained jointly with the diffusion model to extract meaningful conditional features, the decoder is pre-trained as part of the VAE and kept frozen during DiT training, forming an asymmetric training strategy that stabilizes the reconstruction process and improves convergence.
Latent-space diffusion via DiT. The core of AeroDiT is a transformer-based diffusion model (DiT) that iteratively denoises the latent representation. At each reverse timestep, the DiT takes three inputs: (1) the current noisy latent, (2) the condition embedding from the encoder, and (3) a learned timestep embedding.
We adopt a lightweight DiT variant consisting of 6 transformer blocks, each with 8 attention heads and feedforward sublayers. Self-attention is applied over the flattened $H^{\prime}\times W^{\prime}$ latent grid, allowing the model to capture both local structures and long-range dependencies-critical for accurately modeling multiscale turbulence and boundary-layer dynamics. By operating in the compressed latent space, the model achieves strong predictive accuracy while remaining computationally tractable. Compared to pretrained vision models (e.g., ViT or CLIP), which are optimized for natural images, our task-specific encoder provides physics-aware inductive bias and is trained jointly with the diffusion model to ensure consistency across modules.
Decoding and prediction. After iterative denoising in latent space, the resulting tensor is passed through a symmetric convolutional decoder to reconstruct the physical fields at the original resolution $[B,3,H,W]$, corresponding to the predicted pressure and velocity fields.
Together, this architecture enables conditional generation of high-resolution flow fields from geometry and inflow conditions, while ensuring multiscale awareness and physical consistency. The training strategy and hybrid loss functions are detailed in Section III.6.
III.4 Forward Diffusion Process
@@
III.5 Reverse Denoising Process and the Loss
@@
III.6 Physics Informed Loss
The training of AeroDiT is guided by a hybrid loss function that combines data-driven supervision with physics-based constraints. This design enables the model to learn not only from statistical correlations in the training data, but also from the governing laws of fluid mechanics.
The core data-driven term is the mean squared error (MSE) derived in the previous section, to enhance the physical plausibility of the generated flow fields, we incorporate additional loss terms derived from the steady-state Reynolds-Averaged Navier-Stokes (RANS) equations. For incompressible flows, the RANS equations are fundamentally composed of two conservation laws: mass conservation and momentum conservation. These two constraints fully characterize the behavior of the fluid system and are thus used as the basis for our physics-informed loss.
The mass conservation loss penalizes deviations from the divergence-free condition of the velocity field:
$\mathcal{L}_{\text{mass}}=\left\|\nabla\cdot\mathbf{u}\right\|_{2}^{2}=\left\|\frac{\partial u_{x}}{\partial x}+\frac{\partial u_{y}}{\partial y}\right\|_{2}^{2}.$ (10)
The momentum conservation loss measures the residual of the steady incompressible Navier–Stokes equations (per unit mass) by capturing convective transport, pressure gradients, and viscous diffusion. In our implementation, each term is normalized by its mean absolute value to improve numerical stability during training:
$\mathcal{L}_{\text{mom}}$ $=\left\|\frac{(\mathbf{u}\!\cdot\!\nabla)u_{x}}{\sigma_{c}}\;+\;\frac{\partial p}{\partial x}\frac{1}{\sigma_{p}}\;-\;\frac{\mu}{\rho}\frac{\nabla^{2}u_{x}}{\sigma_{v}}\;-\;g_{x}\right\|_{2}^{2}$ (11)
$\quad+\;\left\|\frac{(\mathbf{u}\!\cdot\!\nabla)u_{y}}{\sigma_{c}}\;+\;\frac{\partial p}{\partial y}\frac{1}{\sigma_{p}}\;-\;\frac{\mu}{\rho}\frac{\nabla^{2}u_{y}}{\sigma_{v}}\;-\;g_{y}\right\|_{2}^{2},$
where $\sigma_{c}$, $\sigma_{p}$, and $\sigma_{v}$ denote the mean absolute values of the convective, pressure, and viscous terms, respectively, used for normalization. Here $\nabla^{2}u_{x}$ and $\nabla^{2}u_{y}$ are computed component-wise (vector Laplacian), and $\mathbf{g}=(g_{x},g_{y})$ represents the body force per unit mass.
These components are combined with the MSE loss to form the total training objective:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{MSE}}+\lambda_{\text{phys}}\left(\lambda_{\text{mass}}\mathcal{L}_{\text{mass}}+\lambda_{\text{mom}}\mathcal{L}_{\text{mom}}\right),$ (12)
where $\lambda_{\text{mass}}$ and $\lambda_{\text{mom}}$ control the balance between the two physical terms, and $\lambda_{\text{phys}}$ regulates the overall contribution of physics-based supervision.
These scalar weights play two key roles. First, they address potential scale mismatches between different loss components-particularly between gradients of pressure and velocity fields-by normalizing their influence on training. Second, they provide finer control over the optimization dynamics, enabling us to tune the strength of physical regularization in a principled manner.
To further improve training stability and convergence, we adopt a warm-up strategy for $\lambda_{\text{phys}}$. At the beginning of training, physics-based losses are down-weighted to allow the model to first capture coarse patterns from data. The value of $\lambda_{\text{phys}}$ is then gradually increased to its target level over a fixed number of iterations. This prevents early-stage overfitting to noisy residuals and allows the model to integrate physical constraints more effectively as it learns.
In summary, the hybrid loss formulation equips AeroDiT with both statistical learning capacity and inductive physical bias, enabling more accurate and physically consistent predictions.
III.7 Overall Training
As illustrated in Figure 1, the training pipeline of AeroDiT involves three major components: the condition encoder, the diffusion transformer (DiT), and the function space decoder. The encoder and decoder are pretrained using reconstruction objectives, while the DiT is subsequently trained using the loss function described in Section III.6.
The condition encoder processes airfoil geometry and inflow conditions into latent features. These features are then injected into the DiT either via concatenation or attention mechanisms to condition the generative process.
The DiT module models the conditional distribution $p(\bm{x}\mid\bm{\Psi})$ by denoising noisy flow fields. At each diffusion timestep, the model predicts the added noise, supervised by both MSE and physics-informed losses.
The function space decoder transforms the denoised latent features back into physical fields, such as pressure and velocity. It is responsible for translating abstract features into spatially coherent predictions.
The complete training loop is summarized below.